OpenStudy (blackops2luvr):

What is the height of a cone with a radius of 4 inches and of volume 8 cubic inches? Round your answer to the tenths place. 0.5 inch 1.4 inches 2.0 inches 4.0 inches

OpenStudy (shinalcantara):

\[V = \frac{ 1 }{ 3 } \pi r^2 h\]

OpenStudy (blackops2luvr):

o-0

OpenStudy (shinalcantara):

plug in the given values and you'll find h

OpenStudy (blackops2luvr):

all i wanted was an explanation then an answer on how you got it

OpenStudy (shinalcantara):

(-_-) \[V = \frac{ 1 }{ 3 } \pi r^2 h\] Given: V=8 cubic inches r = 4 inches \[8 = \frac{ 1 }{ 3 } \pi (4)^2 h\] \[h = \frac{ 8(3) }{ \pi 4^2 }\]

OpenStudy (blackops2luvr):

well, h ≈ 0.48 in

OpenStudy (blackops2luvr):

lol

OpenStudy (blackops2luvr):

\(\normalsize\sf\color{black}{can~you~teach~my~how~do~do~that~LaTeX?}\)